In any triangle ABC show that  the following
tan (A+B-C/2)=cot (C)

Grade:11

1 Answers

Nishant Vora IIT Patna
7 years ago
A+ B = 180 – C

A+ B -C = 180 – 2C

A+ B -C / 2 = 180 – 2C / 2 = 90 – C

Hence tan (90 – C) = cot (C)